OpenStudy (anonymous):

(2 pts) A water truck's tank has internal dimensions of 13 meters in length and 1.5 meters in diameter. If the truck's load of pure water is 70 percent of full capacity, how heavy is the load in pounds? Water's density is 62.4 lb/ft3.

OpenStudy (anonymous):

\[\pi*D*L = Volume of tank\] 70% of tank volume is: \[0.7*\pi*D*L\] Density of water in metric units is also: \[1000 kg/m ^{3}\] You now multiply \[(0.7*\pi*D*L) * (1000 kg/m ^{3}) = 42,883 kg\] To get the 'heaviness' or weight of this thing in Newtons (N) --> multiply by gravity acceleration. Weight = \[mass*gravity =(42883 kg)*(9.81 m/s ^{2}) = 420,680 N\] To get in pounds force, you simply 'google' the conversion factor between 1 N and and pounds. That conversion is: 1 lbs = 4.45 N So now just divide that value above by 4.45N and your set in lbs